Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 35     Entry time: 25 Feb 2024, 22:54
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.10 keV; 190 pulses per train @1.1 MHz with roughly 170-200 uJ; stable delivery

Optical laser delivery

  • ns laser for particles visualization

Achievements/Observations

  • Sample injection and data collection for ~11 hours
  • collected data on copper cubes with better hit rates than previous shifts; ~15%

Issues

  • We had laser interlock issues, the magnet in door went to error state and this was fixed by Klaus Witt call around 22:15.
  • data migration issues; it was very slow and DOC did not figure out the cause for this.
ELOG V3.1.4-7c3fd00